# Reasoning-Enhanced Object-Centric Learning for Videos

### v1

@@ 5 Conclusion
We have designed a more reasonable prediction module called STATM, which clearly improved slot-based models in the context of scene understanding and prediction. We demonstrated that reasoning and prediction abilities influence the model’s scene object segmentation and tracking. The more accurate the reasoning and prediction abilities, the stronger the segmentation and tracking of objects. Through a series of experiments, we investigated the influence of memory and spatiotemporal reasoning on the model’s perceptual abilities.Although there still remain many challenges on this topic, the results in this paper illustrate that well-designed deep learning models can mimic human perception. In the future, we will continue to explore more cognitive theories as a basis and continue to work hard to obtain a model that performs well in both perception and prediction.
Appendix A Baselines
For the SAVi model, we chose the official implementation of SAVi with a small CNN as backbone, which is trained and evaluated on downscaled 64*64 frames. We initialized the slots in the first frame using bounding boxes as hints (slot initialization is a simple MLP). Since one of the main contributions of the SAVi model is using optical flow as a self-supervised objectives, we selected optical flow as the target for training the model. During training, we adjusted the batch size to 32, while keeping other settings and parameters the same as SAVi-small.
SAVi++. SAVi++ (Elsayed et al., 2022) has a structure similar to SAVi with a ResNet34 backbone and consists of five main components: encoder, decoder, slot initialization, corrector, and predictor. Unlike SAVi, SAVi++ introduces depth as self-supervised objectives. It also incorporates data augmentation and utilizes a transformer encoder after ResNet34.
During the training of SAVi++, we adjusted the batch size to 32, the number of training steps to 100k, while keeping other parameters consistent with described in SAVi++. For $*{}^{*}$, the hyperparameters are same as the official implementation.
STEVE. STEVE (Singh et al., 2022) is an unsupervised object-centric scene decomposition model. This baseline employs discrete VAE (Im et al., 2017) for encoding and reconstructing input frames $x_{t}$ and generating discrete targets (hard sampling) for the transformer decoder. It uses a similar structure combined with the encoder, corrector, and predictor in SAVi, called the recurrent slot encoder, to decompose input video frames $x_{t}$ into slots. The slot-transformer decoder uses the slots obtained from the recurrent slot encoder to learn to predict the sampling targets from discrete VAE by minimizing the cross-entropy loss. We make no modifications to the official implementation of STEVE.
Appendix B Additional Training Setup
Experiments of STATM-SAVi and STATM-SAVi++. Referring to Section 4.1, we train our models (STATM-SAVi and STATM-SAVi++) for 100k steps with a batch size of 32 using Adam (Kingma & Ba, 2014). Same as SAVi++ (Elsayed et al., 2022), we linearly increase the learning rate for 5000 steps to 0.0002 (starting from 0) and then decay the learning rate with a Cosine schedule (Loshchilov & Hutter, 2016). We split each video into sub-sequences of 6 frames to train the model (In the ablation experiment studying the impact of training stage buffer size on the model, sub-sequences are set to 4/6/12, referring to Section 4.2). In the initialization of slots for the first frame, bounding boxes are utilized as contextual cues. For MOVi-A, B, and C datasets, the number of slots is set to 11. In the case of datasets MOVi-D and E, the number of slots is set to 24. We use 1 iteration per frame for the Slot Attention (Locatello et al., 2020) module. All other parameters and settings for each model remain consistent with their respective baselines. We implement models in JAX using the Flax neural network library, unless stated otherwise. We train all models on 8*4090 GPUs with 24GB memory each (We also trained the model on A100 with 80GB memory or 3090 with 24GB memory. However, for the sake of convenience in time calculations, we standardized the timing to the specifications of the 4090). STATM-SAVi-small takes approximately 8hs for the MOVi-A, B and C datasets, and 9hs for the MOVi-D and E datasets. STATM-SAVi++ takes approximately 9hs for the MOVi-A, B and C datasets, and 17hs for the MOVi-D and E datasets.
Experiments of $*{}^{*}$. Further, we modify the training steps to 500k and change the batch size to 64 to train the STATM-SAVi++ model. Additionally, for the Slot Attention on the MOVi-E dataset, we adjust the number of iterations to 2 per frame. All other parameters and settings remain consistent with the SVAi++ paper (Elsayed et al., 2022). We train the model on the MOVi-A, B and C datasets using 8*4090 GPUs with 24GB memory each, which takes about 4 days. For the MOVi-D and E datasets, we used 8*H800 GPUs with 80GB memory each to train the model, and the training process takes about 8 days.
Experiments of STATM-STEVE. Due to the necessity of memory in STATM, we modify the training sub-sequence length to 3/6/24 (corresponding to batch sizes of 24/12/8). All other parameters and settings remain consistent with the STEVE paper (Singh et al., 2022). The model is trained on 2*A100 with 80GB memory each, and 500 epochs require approximately 6 days. We implement STATM-STEVE in PyTorch (Paszke et al., 2019).
Appendix C Additional Parameter
STATM-SAVi-Small indeed has a parameter increase of approximately 66K compared to SAVi-Small, which is notably smaller than the parameter increase seen in SAVi-Large compared to SAVi-Small (around 21378K parameters). Moreover, our STATM-SAVi-Small model, trained for 100k steps with a batch size of 32, performs similarly to the official SAVi-Large model, trained for 500k steps with a batch size of 64. This further highlights the reasonableness and superiority of our designed prediction module.
Table 3: Comparison of the parameter number for different models.
@@ Appendix D Additional Experimental Results
Additional Segmentation Results. In order to better assess our model, we conducted an evaluation using the first 6 frames of the videos. Detailed results can be found in Table 4. Referring to Table 1 in the main text, we can observe the following trends: on simple datasets, the decline in our model’s object segmentation and tracking capabilities over extended time sequences is comparable to that of the baseline model. However, on complex datasets like MOVi-C, D, and E, the decrease in our model’s performance is significantly less than that of the baseline model. This indicates that the STATM is more suitable for handling object segmentation and tracking tasks in longer-time sequences and complex environments. This finding further validates the effectiveness of our STATM model.
Table 4: Segmentation results on the first 6 frames of the MOVi dataset.
@@ A B C D E A B C D E
Additional Comparison of STEVE Results. STEVE (Singh et al., 2022) is a unsupervised baselines which is state-of-the-art in object-centric video scene decomposition. It includes a structure similar to SAVi, consisting of an encoder, corrector, and predictor, referred to as the recurrent slot encoder. Therefore, we modified the STEVE model using the STATM module and conducted a brief evaluation of the modified model on the MOVi-D and E datasets. (Note that the STATM structure in STATM-STEVE differs slightly from that in STATM-SAVi. For details, please refer to Appendix F).
From the FG-ARI scores in Table 5, it can be observed that the STATM module improves the segmentation performance of the STEVE on MOVi-D and E datasets. However, STEVE mainly focuses on scene segmentation and does not report mIoU, we don’t compare mIoU. During the experiments, we found that the STATM-STEVE model is sensitive to the MSE loss in discrete VAE (Im et al., 2017), and the addition of STATM slightly increases the difficulty of fitting of the model, which becomes more pronounced with an increase in the buffer size of STATM module. Therefore, for a better improvement and testing of the impact of STATM on STEVE, a two-stage training approach can be attempted, where a discrete VAE is first trained, followed by the training of STATM and other modules.
Table 5: Segmentation results on the MOVi-D and E datasets.
Model MOVi-D MOVi-E
STEVE 47.67 52.15
STATM-STEVE $\bm{51.73}$ $\bm{55.78}$
Additional Qualitative Results. We show more qualitative results on longer time series in Figure 6 to Figure 10.
@@
Visualization of the Attention Map. To analyze slots and better illustrate the relationship between objects and slots, we visualized the attention map of the Slot Attention (corrector) in Figure 11 to Figure 15.
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 6: Qualitative results of our model compared to SAVi++ on the MOVi-A dataset.
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 7: Qualitative results of our model compared to SAVi++ on the MOVi-B dataset.
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 8: Qualitative results of our model compared to SAVi++ on the MOVi-C dataset.
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 9: Qualitative results of our model compared to SAVi++ on the MOVi-D dataset.
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Video Refer to caption
G.T. Refer to caption
SAVi++ Refer to caption
Ours++ Refer to caption
Figure 10: Qualitative results of our model compared to SAVi++ on the MOVi-E dataset.
Figure 11: Attention map visualization on the MOVi-A dataset.
Figure 12: Attention map visualization on the MOVi-B dataset.
Figure 13: Attention map visualization on the MOVi-C dataset.
Figure 14: Attention map visualization on the MOVi-D dataset.
Figure 15: Attention map visualization on the MOVi-E dataset.
Additional Generalization Results. We selected the models trained with a batch size of 32 and 100k training steps to assess its generalization. The test set utilized the default test split of MOVi-C, D and E dataset, featuring scenes exclusively consist of held-out objects and background images to evaluate generalization. The results are presented in Table 6.
Table 6: OOD generalization results of STATM-SAVi and STATM-SAVi++ on the MOVi dataset.
[…]
C D E C D E
STATM-SAVi-S (IID) 34.0 17.0 9.0 57.7 40.9 36.9
STATM-SAVi-S (OOD) 33.2 17.0 8.2 59.7 43.4 35.9
STATM-SAVi++ (IID) 49.5 50.1 27.9 77.7 85.8 85.0
STATM-SAVi++ (OOD) 48.7 50.3 27.4 78.9 86.4 85.8
Appendix E Additional Ablation Experiment of Memory Module
Training with unlimited buffer length and testing with limited buffer length. To better assess the impact of the buffer on the model, we trained the model using video sub-sequences of 12 frames, as shown in Table 7 and 8. We observed that: 1) On relatively simple datasets like MOVi-A, B, C, and D dataset, increasing the amount of training data with additional information would enhance the model’s segmentation capabilities. 2) Training with a buffer size of 12 results in a decrease in mIoU. SAVi++ augments the number of samples by utilizing sub-sequences of 6 frames. However, in this case, sub-sequences of 12 frames are employed, leading to a reduction in both the number of samples and hints (each sample contains hints in the first frame, so the reduction in the number of samples also reduces the hints). The decrease in both samples and hints may impact the model’s ability to separate foreground and background, consequently causing a decline in mIoU. 3) On the MOVi-E dataset, increasing the number of training frames resulted in a decrease in the model’s tracking and segmentation capabilities. This could be attributed to the limitations in the ability of the upstream modules to effectively extract image features. The findings from the SAVi and SAVi++, which used more powerful encoders and data augmentation to improve segmentation performance on MOVi-E, support this observation. Therefore, exploring the design of a more robust encoder and refining the corrector and guidance modules may yield unexpected improvements. We plan to further investigate this direction in future research.
Table 7: Evaluation on all video frames of the model trained using 12 frames (B represents the size of the buffer during the testing phase).
@@
Table 8: Evaluation on the first 6 video frames of the models trained by 6 frames and 12 frames (T represents the number of frames used for training model, B represents the size of the buffer during the testing phase).
@@
Training with limited buffer length and testing with limited buffer length. We intentionally limited the buffer size during both the training and testing phases, and the model evaluation results are presented in Table 9. Remarkably, we found that the model trained with a smaller buffer (equivalent to restricted vision) experienced less impact from the buffer during the testing phase. For instance, consider a model trained with a buffer size of 2. When tested with a reduced buffer size on the MOVi-E dataset, the model experienced an approximately 8% decrease in FG-ARI (from 23.6% to 15.5%). On the other hand, when testing with a reduced buffer size on the MOVi-E dataset, a model trained with a buffer size of 6 exhibited a FG-ARI decrease of about 30% (from 36.8% to 6.8%). This has intriguing implications for the fusion of deep learning and cognitive science. However, it’s important to note that real human learning and cognitive processes are likely more complex and influenced by various factors. This study provides a theoretical framework, but further theoretical substantiation and experimental validation are still needed.
Table 9: Evaluation result of the model trained limited buffer length (T represents the size of the buffer during the training phase, B represents the size of the buffer during the testing phase).

### v2

@@ 5. Conclusion
In the real world, all objects follow the laws of physics. Intuitive physics serves as the bridge and connection through which humans comprehend the world. Our research aims to construct an object-centric system for object perception, learning of physics, and reasoning to explore whether deep learning models can learn physical concepts like humans, and use these learned physical laws to make inferences and predictions about the future motion of objects. We have designed a more reasonable prediction module called STATM, which clearly improved slot-based models in the context of scene understanding and prediction. We demonstrated that reasoning and prediction abilities influence each other. Through a series of experiments, we have shown the advantages of our model in tasks such as perception, prediction, and VQA. We also explore the impact of different spatiotemporal attention and fusion methods, and memory buffer on model perception and prediction. Although many challenges still remain in this field, the results presented in this paper illustrate that well-designed deep learning models can mimic human perception and prediction. In the future, we hope to implement joint training of our perception and prediction models, along with real-time perception and prediction, and validate the model in real-world scenarios.
Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 92270118, No. 62276269), the Beijing Natural Science Foundation (No. 1232009), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0620103). H. Sun and Y. Liu would like to thank the support from the Fundamental Research Funds for the Central Universities (No. 202230265, No. E2EG2202X2).
Appendix
This supplementary material file provides the appendix section to the main article.
Table S1. Comparison of the parameter number for different models.
@@ Acknowledgements.
Appendix A Baselines
For the SAVi model in perception, we chose the official implementation of SAVi with a small CNN as backbone, which is trained and evaluated on downscaled 64*64 frames. We initialized the slots in the first frame using bounding boxes as hints (slot initialization is a simple MLP). We selected optical flow as the target for training the model. During training, we adjusted the batch size to 32, while keeping other settings and parameters the same as SAVi-small.
SAVi++. SAVi++ (Elsayed et al., 2022) has a structure similar to SAVi. During the training of SAVi++, we adjusted the batch size to 32, the number of training steps to 100k, while keeping other parameters consistent with described in SAVi++. For SAVi++, the hyperparameters are same as the official implementation.
STEVE. STEVE (Singh et al., 2022) is an unsupervised object-centric scene decomposition model. This baseline employs discrete VAE (Im et al., 2017) for encoding and reconstructing input frames $x_{t}$ and generating discrete targets for the transformer decoder. It uses a similar structure combined with the encoder, corrector, and predictor in SAVi, called the recurrent slot encoder, to decompose input video frames $x_{t}$ into slots. The slot-transformer decoder uses the slots obtained from the recurrent slot encoder to learn to predict the sampling targets from discrete VAE by minimizing the cross-entropy loss. We make no modifications to the official implementation of STEVE.
SlotFormer. SlotFormer (Wu et al., 2023) is a transformer-based framework for object-centric visual simulation. It leverages slots extracted by upstream modules like SAVi to train a slot-based transformer encoder model for prediction purposes. Additionally, it utilizes results from rollout simulations to train Aloe (Ding et al., 2021b) for Visual Question Answering (VQA) tasks. We employ an unsupervised approach to train SAVi-small on CLEVRER for slot extraction, with all other configurations consistent with the official settings.
SAVi-SlotFormer. To evaluate whether enhancements to the predictor can alter the performance of SAVi, we replaced the predictor in SAVi with SlotFormer, which is currently the best-performing model in object-centric prediction, featuring a memory buffer and transformer encoders predictor. All other settings are consistent with SAVi.
G-SWM. G-SWM (Lin et al., 2020) is an unsupervised, object-centric predictive model that calculates foreground and background distributions through two separate modules, subsequently rendering and combining these results for dynamic prediction. Object interactions and occlusions are managed through a simple graph neural network. The official implementation has been trained using CLEVRER, yielding results comparable to those obtained with SlotFormer (Wu et al., 2023).
SAVi-dyn. In SlotFormer (Wu et al., 2023), Wu et al. enhanced prediction capabilities by replacing the Transformer predictor in SAVi with a Transformer-LSTM module in PARTS (Zoran et al., 2021). We trained SAVi-dyn using the same setup as theirs.
DCL. DCL (Chen et al., 2021b) utilizes a trajectory extractor to monitor each object over time, representing it as a latent, object-centric feature vector. Building on this foundational representation, DCL employs graph networks to learn and approximate the dynamic interactions among objects.
Aloe. Aloe (Ding et al., 2021b) trains transformers using slots for prediction. To enable a direct comparison with models such as SlotFormer, we use the Aloe model as re-implemented in the SlotFormer paper, ensuring all hyperparameters and settings are kept consistent with those described in the paper.
Appendix B Additional Training Setup
Experiments of STATM-SAVi and STATM-SAVi++. Referring to Section 4.1, we train our models (STATM-SAVi and STATM-SAVi++) for 100k steps with a batch size of 32 using Adam (Kingma and Ba, 2014). Same as SAVi++ (Elsayed et al., 2022), we linearly increase the learning rate for 5000 steps to 0.0002 (starting from 0) and then decay the learning rate with a Cosine schedule (Loshchilov and Hutter, 2016). We split each video into sub-sequences of 6 frames to train the model (In the ablation study studying the impact of training stage buffer size on the model, sub-sequences are set to 4/6/12, referring to Section 4.3). In the initialization of slots for the first frame, bounding boxes are utilized as contextual cues. For MOVi-A, B, and C datasets, the number of slots is set to 11. In the case of datasets MOVi-D and E, the number of slots is set to 24. We use 1 iteration per frame for the Slot Attention (Locatello et al., 2020) module. All other parameters and settings for each model remain consistent with their respective baselines. We implement models in JAX using the Flax neural network library, unless stated otherwise.
Experiments of STATM-SAVi++. Further, we modify the training steps to 500k and change the batch size to 64 to train the STATM-SAVi++ model. Additionally, for the Slot Attention on the MOVi-E dataset, we adjust the number of iterations to 2 per frame. All other parameters and settings remain consistent with the SVAi++ (Elsayed et al., 2022).
Experiments of STATM-STEVE. Due to the necessity of memory in STATM, we modify the training subsequence length to 3/6/24 (corresponding to batch sizes of 24/12/8). During the experiments, we found that the STATM-STEVE model is sensitive to the MSE loss in discrete VAE (Im et al., 2017), and the addition of STATM slightly increases the difficulty of fitting of the model, which becomes more pronounced with an increase in the buffer size of STATM module. Therefore, for a better improvement and testing of the impact of STATM on STEVE, a two-stage training approach can be attempted, where a discrete VAE is first trained, followed by the training of STATM and other modules. All other parameters and settings remain consistent with the STEVE (Singh et al., 2022). We implement STATM-STEVE in PyTorch.
Experiments of Prediction and VQA. We train our models (STATM-SAVi) for 400k steps with a batch size of 64 on the CLEVRER dataset to extract slots. The number of slots is set to 7, with a learning rate of 0.0001. For prediction, we subsample the video by a factor of 2 to train STATM, conducting approximately 500k training steps with a batch size of 64 and a learning rate of 0.0002. We use rollout slots to train Aloe, targeting around 300k steps with a learning rate of 0.0001 and a batch size of 128. We use the Adam optimizer and apply the same warm-up and decay learning rate schedule for the first 2.5% of the total training steps.
Appendix C Additional Parameter
STATM-SAVi-Small has a parameter increase of approximately 66K compared to SAVi-Small, which is notably smaller than the parameter increase seen in SAVi-Large compared to SAVi-Small (around 21378K parameters). Moreover, our STATM-SAVi-Small model, trained for 100k steps with a batch size of 32, performs similarly to the official SAVi-Large model, trained for 500k steps with a batch size of 64. This further highlights the reasonableness and superiority of our designed prediction module.
@@ Appendix D Additional Experimental Results
Table S2. Segmentation results on the first 6 frames of the MOVi dataset.
@@ A B C D E A B C D E
Additional Segmentation Results. In order to better assess our model, we conducted an evaluation using the first 6 frames of the videos. Detailed results can be found in Table S2. Referring to Table 1 in the main text, we can observe the following trends: on simple datasets, the decline in our model’s object segmentation and tracking capabilities over extended time sequences is comparable to that of the baseline model. However, on complex datasets like MOVi-C, D, and E, the decrease in our model’s performance is significantly less than that of the baseline model. This indicates that the STATM is more suitable for handling object segmentation and tracking tasks in longer-time sequences and complex environments. This finding further validates the effectiveness of our STATM model.
Table S3. Accuracy on different questions and average results on CLEVRER.
Model Descriptive Explanatory Predictive Counterfactual Average
per opt. per ques. per opt. per ques. per opt. per ques.
DCL 90.70 89.58 82.82 90.52 82.03 80.38 46.52 75.52
VRDP 93.40 96.30 91.94 95.68 91.35 94.83 84.29 90.24
SlotFormer 95.17 98.04 94.79 96.50 93.29 90.63 73.78 89.26
STATM(Ours) 95.22 98.15 95.04 96.62 93.63 90.57 73.90 89.44
@@
Additional Qualitative Results. We show more qualitative results on longer time series in Figure S1 to Figure S5. Meanwhile, To analyze slots and better illustrate the relationship between objects and slots, we visualized the attention map of the Slot Attention (corrector) in Figure S6 to Figure S10.
Additional VQA Results. Detailed results about all questions on CLEVRER are presented in Table S3.
Appendix E Additional Ablation Study
Table S4. Evaluation on all video frames of the model trained using 12 frames (B represents the size of the buffer during the testing phase).
@@
Table S5. Evaluation on the first 6 video frames of the models trained by 6 frames and 12 frames (T represents the number of frames used for training model, B represents the size of the buffer during the testing phase).
@@
Table S6. Evaluation result of the model trained limited buffer length (T represents the size of the buffer during the training phase, B represents the size of the buffer during the testing phase).
@@ A E A E A E A E
Table S7. Accuracy on different questions and average results of different buffer sizes on CLEVRER.
Model Descriptive Explanatory Predictive Counterfactual Average
per opt. per ques. per opt. per ques. per opt. per ques.
STATM(24) 95.22 98.15 95.04 96.62 93.63 90.57 73.90 89.44
STATM(32) 95.34 98.23 95.30 96.41 93.09 90.91 74.69 89.61
STATM(48) 95.17 98.14 95.14 96.07 92.67 91.01 74.51 89.37
STATM(128) 95.15 97.84 94.15 96.21 92.90 90.50 73.44 88.91
Training with unlimited buffer length and testing with limited buffer length. To better assess the impact of the buffer on the model, we trained the model using video sub-sequences of 12 frames, as shown in Table S4 and Table S5. We observed that: 1) On relatively simple datasets like MOVi-A, B, C, and D dataset, increasing the amount of training data with additional information would enhance the model’s segmentation capabilities. 2) Training with a buffer size of 12 results in a decrease in mIoU. SAVi++ augments the number of samples by utilizing sub-sequences of 6 frames. However, in this case, sub-sequences of 12 frames are employed, leading to a reduction in both the number of samples and hints. The decrease in both samples and hints may impact the model’s ability to separate foreground and background, consequently causing a decline in mIoU. 3) On the MOVi-E dataset, increasing the number of training frames resulted in a decrease in the model’s tracking and segmentation capabilities. This could be attributed to the limitations in the ability of the upstream modules to effectively extract image features. The findings from the SAVi and SAVi++, which used more powerful encoders and data augmentation to improve segmentation performance on MOVi-E, support this observation. Therefore, exploring the design of a more robust encoder and refining the corrector and guidance modules may yield unexpected improvements. We plan to further investigate this direction in future research.
Training with limited buffer length and testing with limited buffer length. We intentionally limited the buffer size during both the training and testing phases, and the model evaluation results are presented in Table S6. Remarkably, we found that the model trained with a smaller buffer experienced less impact from the buffer during the testing phase. For instance, consider a model trained with a buffer size of 2. When tested with a reduced buffer size on the MOVi-E dataset, the model experienced an approximately 8% decrease in FG-ARI (from 23.6% to 15.5%). On the other hand, when testing with a reduced buffer size on the MOVi-E dataset, a model trained with a buffer size of 6 exhibited a FG-ARI decrease of about 30% (from 36.8% to 6.8%). This has intriguing implications for the fusion of deep learning and cognitive science. However, it’s important to note that real human learning and cognitive processes are likely more complex and influenced by various factors. This study provides a theoretical framework, but further theoretical substantiation and experimental validation are still needed.
VQA results of different buffer size. Detailed results for all questions on CLEVRER using different memory buffer sizes are presented in Table S7.